OpenStudy (anonymous):

What is simplified version of -n+3/n-2 + -3n-7/n-2

OpenStudy (jhannybean):

\[\sf \frac{-n+3}{n-2}+\frac{-3n-7}{n-2} \implies \frac{(-n+3)-3n-7}{n-2}\]

OpenStudy (steve816):